### v1

@@ 0.4.3 Adversarial Training on Semantic Subspaces
    We use $N_{base}$ as the baseline network from the previous experiments for comparison with networks trained for robustness.
@@
    Data augmentation. We obtain three augmented datasets $\mathcal{Y}\cup\mathcal{A}^{pos}_{char}$, $\mathcal{Y}\cup\mathcal{A}^{pos}_{word}$ and $\mathcal{Y}\cup\mathcal{A}^{pos}_{pj}$. Note that $\mathcal{A}^{pos}_{char}$, $\mathcal{A}^{pos}_{word}$ and $\mathcal{A}^{pos}_{pj}$ are in correspondence with the sets defined in Section 0.4.2. The subscripts char and word denote the kind of perturbation, for which the types are detailed in Tables 3 and 4, while the subscript pj refers to the sentence level perturbations generated with Polyjuice. We train the baseline architecture, using the standard stochastic gradient descent and cross entropy loss, on the augmented datasets, and obtain DNNs $N_{char-aug}$, $N_{word-aug}$ and $N_{pj-aug}$.
  3. 3.
    Customised PGD adversarial training. As defined in Section 0.3.4, we vary the parameter $\mathbb{H}$. In particular, we take for $\mathbb{H}$ different choices of geometric and semantic subspaces we have defined so far. We denote a network trained with PGD algorithm on a subspace $\mathbb{H}^{*}_{name}$ as $N_{name}$. For example, taking $\mathbb{H}=\mathbb{H}_{sh}$ we get $N_{sh}$ as a result of adversarially training the benchmark architecture on $\mathbb{H}_{sh}$. See Table 13 for full listing of the networks we obtain in this way.
We call DNNs of second and third type robustly trained networks. We keep the geometric and semantic subspaces from the previous experiments (shown in Table 9) to compare how training affects their verifiability. Following the same evaluation methodology of experiments as in Sections 0.4.1 and 0.4.2, we use the verifiers ERAN and Marabou to measure verifiability of the subspaces. Table 13 reports accuracy of the robustly trained networks, while the verification results are presented in Tables 14 and 15. From Table 13 we can see that networks trained with data augmentation achieve similar nominal accuracy to networks trained with adversarial training. However, the most prominent difference is exposed in Tables 14 and 15: adversarial training effectively improves the verifiability of the networks, while data augmentation actually decreases it.
Specifically, the adversarially trained networks trained on semantic subspaces ($N_{char}$, $N_{word}$, $N_{pj}$) achieved high verifiability, reaching up to $45.87\%$ for R-U-A-Robot and up to $83.48\%$ for the Medical dataset. This constitutes a significant improvement of the verifiability results compared to $N_{base}$. Looking at nuances, there does not seem to be a single winner subspace when it comes to adversarial training, and indeed in some cases $\mathbb{H}^{*}_{\epsilon=0.05}$ wins over more precise subspaces. All of the subspaces in Table 9 have very similar volume, which accounts for improved performance across all experiments. The particular peaks in performance then come down to particularities of a specific semantic attack that was used while training. For example, the best performing networks are those trained with Polyjuice attack, the strongest form of attack in our range. Thus, if the kind of attack is known in advance, the precision of hyper-rectangles can be further tuned.
Model Dataset Adversarial training Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
$N_{char-aug}$
$\mathcal{Y}\cup\mathcal{A}^{pos}_{char}$
No
0.9562 ± 0.0026
0.9320 ± 0.0035
0.9908 ± 0.0006
0.9346 ± 0.0030
$N_{word-aug}$
$\mathcal{Y}\cup\mathcal{A}^{pos}_{word}$
No
0.9857 ± 0.0006
0.9459 ± 0.0036
$N_{pj-aug}$
$\mathcal{Y}\cup\mathcal{A}^{pos}_{pj}$
No
0.9819 ± 0.0009
0.9319 ± 0.0039
$N_{char}$
$\mathcal{Y}$
Yes
0.9326 ± 0.0019
0.9251 ± 0.0038
0.9627 ± 0.0005
0.9509 ± 0.0016
$N_{word}$
$\mathcal{Y}$
Yes
0.9368 ± 0.0016
0.9237 ± 0.0029
$N_{pj}$
$\mathcal{Y}$
Yes
0.9505 ± 0.0019
0.9349 ± 0.0032
$N_{\epsilon=0.05}$
$\mathcal{Y}$
Yes
0.9401 ± 0.0017
0.9224 ± 0.0019
0.9605 ± 0.0009
0.9504 ± 0.0024
Table 13: Accuracy of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets. $\mathcal{Y}$ stands for R-U-A-Robot or Medical depending on the column.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{char-aug}$ 0.00 0.24 0.00 0.51 1.38 1.09 0.35 1.06 -
$N_{word-aug}$ 0.00 0.24 0.00 0.42 0.31 0.57 0.25 0.92 -
$N_{char}$ 0.00 8.97 4.43 4.81 9.86 11.3 6.91 8.51 -
$N_{word}$ 0.04 10.75 4.05 4.36 8.60 9.52 6.81 7.45 -
$N_{\epsilon=0.05}$ 0.12 10.16 4.18 4.04 8.91 10.17 6.52 7.36 -
Medical $N_{char-aug}$ 0.00 - 7.59 5.28 12.84 11.05 7.92 7.40 26.97
$N_{pj-aug}$ 0.00 - 10.31 8.49 15.67 14.90 9.18 10.58 28.59
$N_{char}$ 5.28 - 50.12 49.78 53.99 57.76 48.02 52.07 55.44
$N_{pj}$ 2.83 - 47.11 46.14 52.12 56.14 44.59 48.27 57.36
$N_{\epsilon=0.05}$ 8.68 - 51.60 50.31 55.67 58.52 50.10 53.65 59.76
Table 14: Verifiability of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for ERAN verifier.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{char-aug}$ 0.72 13.90 8.49 7.92 13.67 15.50 9.56 11.88 -
$N_{word-aug}$ 0.24 11.30 3.87 4.05 8.27 8.84 5.71 7.72 -
$N_{char}$ 7.37 41.93 30.41 30.23 38.20 45.87 32.74 36.62 -
$N_{word}$ 12.17 45.12 25.82 25.39 33.85 37.45 26.87 30.99 -
$N_{\epsilon=0.05}$ 18.46 41.93 21.99 20.32 28.13 32.83 23.52 26.74 -
Medical $N_{char-aug}$ 1.14 - 37.05 35.29 41.50 42.47 34.89 37.94 49.65
$N_{pj-aug}$ 5.77 - 39.00 38.66 42.28 44.22 37.29 39.03 38.22
$N_{char}$ 51.70 - 77.59 77.25 77.50 77.98 77.92 78.67 76.58
$N_{pj}$ 57.45 - 81.94 81.47 82.31 83.48 82.47 82.72 82.24
$N_{\epsilon=0.05}$ 62.57 - 79.32 78.57 78.70 80.21 79.40 80.76 66.22
Table 15: Verifiability of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets, for a selection of semantic subspaces; for Marabou verifier.
$N_{sh}$
Yes
0.9339 ± 0.0022
0.9296 ± 0.0013
0.9614 ± 0.0012
0.9429 ± 0.0026
$N_{50}$
Yes
0.9432 ± 0.0014
0.9349 ± 0.0019
0.9556 ± 0.0020
0.9515 ± 0.0012
$N_{100}$
Yes
0.9488 ± 0.0004
0.9418 ± 0.0024
0.9571 ± 0.0011
0.9547 ± 0.0016
$N_{200}$
Yes
0.9509 ± 0.0009
0.9445 ± 0.0014
0.9585 ± 0.0005
0.9543 ± 0.0010
$N_{250}$
Yes
0.9522 ± 0.0008
0.9422 ± 0.0023
0.9607 ± 0.0013
0.9538 ± 0.0022
$N_{\epsilon=0.005}$
Yes
0.9348 ± 0.0021
0.9159 ± 0.0007
0.9624 ± 0.0004
0.9513 ± 0.0009
Table 16: Accuracy of the robustly (adversarially) trained DNNs on the R-U-A-Robot and the Medical datasets.
Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}^{*}_{50}}$ $\boldsymbol{\mathbb{H}^{*}_{100}}$ $\boldsymbol{\mathbb{H}^{*}_{200}}$ $\boldsymbol{\mathbb{H}^{*}_{250}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.005}}$
RUAR $N_{sh}$ 0.00 0.00 1.33 0.52 0.41 88.62
$N_{50}$ 0.00 0.00 0.00 0.00 0.41 90.02
$N_{100}$ 0.00 0.00 0.00 0.00 0.41 92.74
$N_{200}$ 0.00 0.00 0.00 0.00 0.08 93.54
$N_{250}$ 0.00 0.00 0.00 0.00 0.33 93.86
$N_{\epsilon=0.005}$ 0.00 0.00 0.00 0.00 0.33 98.22
Medical $N_{sh}$ 0.00 0.00 0.00 2.50 4.40 97.47
$N_{50}$ 0.00 0.00 1.08 3.60 6.00 98.79
$N_{100}$ 0.00 0.00 1.08 3.00 5.04 99.09
$N_{200}$ 0.00 0.00 1.08 2.90 4.96 99.05
$N_{250}$ 0.00 0.00 0.00 2.90 4.40 98.73
$N_{\epsilon=0.005}$ 0.00 0.00 0.00 2.30 4.32 99.60
Table 17: Verifiability of the robustly (adversarially) trained DNNs on the R-U-A-Robot and the Medical datasets, for a selection of geometric subspaces; for ERAN verifier.
As a final note, we report results from robust training using the subspaces of Experiment 0.4.1 (Table 6). Table 16 reports the accuracy and the details of the robustly trained networks on those subspaces, while the verification results are presented in Table 17. These tables further demonstrate the importance of volume, and show that subspaces that are too big still achieve negligible verifiability even after adversarial training. Generalisability of the shapes used in Tables 13 - 17 remains the same, see Tables 812.
0.5 NLP Case Studies
The purpose of this section is two-fold:
  • Firstly, the case studies we present here apply the NLP Verification Piplene set out in Section 0.2.5 using more modern NLP tools. Notably, in this section we try different LLMs to embed sentences and replace Polyjuice with the LLM vicuna-13b [^8] Using the following API: https://replicate.com/replicate/vicuna-13b/api., a state-of-the-art open source chatbot trained by fine-tuning LLaMA [115] on user-shared conversations collected from ShareGPT  [^9] https://sharegpt.com/. For further details, please refer to Section 0.3.1.
    To support the work with varying different components of the NLP Verification pipeline, we released the tool ANTONIO [21] that implements different choices for each pipeline component, as shown in Figure 4.
    Figure 4: Tool ANTONIO that implements a modular approach to the NLP verification pipeline introduced in this paper.
  • And secondly, and perhaps more fundamentally, the purpose of this section is to draw attention to the importance of the purely NLP parts of the pipeline, namely the parts that have to do with generation, processing, perturbing, and embedding sentences. This part is largely ignored in NLP verification papers. We show in this section that taking NLP decisions for granted may compromise the practical value of the NLP verification pipelines.
    To show what we mean by these let us show some examples:
    Example 3 (Failure of the Embedding Function; Falsifiable Subspaces).
    As an example, imagine a scenario when a DNN was verified on subpaces of a class $c_{j}$. Then it is used on new, yet unseen sentences. Imagine also that the implicit assumption that the embedding function embeds only semantically similar sentences close in the embedding space, fails. Then, the new sentences that should belong to a class $c_{i}$ may be by mistake embedded into the verified subspace (of class $c_{j}$). It will be wrong to assure the user that the verified subspace guarantees that these new sentences belong to $c_{j}$ (even if the DNN is guaranteed to classify them as $c_{j}$)! In fact we will say that new sentences of class $c_{i}$ that fall inside the verified subspace of class $c_{j}$ falsify the verified subspace. Omitting to acknowledge and report the problem of falsifiable verified subspaces may have different implications, depending on the usage scenarios. If the verified subspace of class $c_{j}$ is used to recognise and censor sensitive (‘dangerous’) sentences, new sentences of class $c_{i}$ that fall inside of the verified subspace will be wrongly censored; which in turn may make interaction with the chatbot impractical. But if the verified subspace of class $c_{j}$ certifies safe sentences, then potentially dangerous sentences of class $c_{i}$ could be wrongly admitted as verifiably safe. Note that this problem can be related to the well-known problem of false positives and false negatives in machine learning: the new sentences of class $c_{i}$ that fall within the verified subpspace of class $c_{j}$ at the same time will be false positives or negatives for that DNN (depending whether $c_{j}$ is a positive or negative class).
    Example 4 (Failure of the Sentence Perturbation Algorithm).
    Another assumption that most NLP verification papers make is that we can algorithmically generate sentence perturbations in a way that is guaranteed to retain their original semantic meaning. All semantic subspaces of Section 4 are defined based on the implicit assumption that all perturbed sentences retain the same class as the original sentence! But if this assumption fails, we will once again end up drawing falsifiable semantic subspaces around embeddings of sentences belonging to different classes, with similar consequences as in the previous example.
As a consequence of the embedding gap, we obtain a seemingly paradoxical situation, when in principle, the same semantic subspace can be both formally verified and empirically falsified! Formal verification ensures that all sentences embedded within the semantic subspace will be classified identically by the given DNN; but empirical falsification of the semantic subspace comes from appealing to the semantic meaning of the embedded sentences, – something that only human can define exactly and verification methodology does not directly define.
In the light of this limitation, How can we measure and improve the quality of the purely NLP components of the pipeline, in a way that ensures that the NLP verification pipeline is sufficiently reliable and practically usable? – is the main question that we answer in this section.
0.5.1 Practicalities of NLP Verification for Critically Safe Applications
Role of false positives and false negatives.
Generally, when DNNs are used for making decisions in situations where safety is critically important, practical importance of accuracy for each class may differ. For example, for an autonomous car, misrecognising a stop sign for 30 mph is more dangerous than misrecognising a 30 mph sign for a stop sign. Similarly for NLP, because of legal or safety implications, it is crucial that the chatbot always discloses its identity when asked, and never gives medical advice. We want to avoid false negatives altogether, i.e. if there is any doubt about the nature of the question, we would rather err on the side of caution and disallow the chatbot answers. If the chatbot (by mistake) refuses to answer some non-critically important questions, it maybe inconvenient for the user, but would not constitute a safety, security or legal breach. Thus, false positives maybe tolerated.
@@ 0.5.1 Practicalities of NLP Verification for Critically Safe Applications
    Firstly, if we use DNN on its own, without verification, we may want to report precision and recall  [^10] Recall that precision is defined as true positivestrue positives+false $\frac{\text{true positives}}{\text{true positives}+\text{false negatives}}$ and recall is defined as true positivestrue positives+false $\frac{\text{true positives}}{\text{true positives}+\text{false positives}}$. in addition to accuracy. Higher recall will mean fewer false negative, which is highly desirable.
@@
    In the literature and in this paper, it is assumed that verified DNNs serve as filters that allow the larger system to use machine learning in a safe manner. Verification covers only subspaces that characterise the critically important class or classes. I.e.: DNN verification guarantees that all meaning-preserving perturbations of a stop sign, or “Are you a Robot?” sentence will be classified by the given DNN in the same way as the original unperturbed sample. In practice, only DNN outputs on these safe regions are trusted, and DNN output bears no verification guarantees for all other inputs. This approach completely eliminates the effect of any false negatives. But the verified subspaces may still contain false positives (cf. Example 3) and in this sense be “falsifiable”, which may not be pragmatically desirable (but should not constitute a security threat).
Bearing these considerations in mind, throughout this section we report precision and recall of models, in addition to accuracy, to see whether one or another class is particularly challenging to recognise. Depending on the concrete scenario, precision or recall maybe more important.
We also introduce the falsifiability metric, that should be used together with verifiability and generalisability metrics in all NLP benchmarks.
0.5.2 Robustness of Large Language Models
We start first with gaining an intuition about robustness of modern LLMs. We will prompt an LLM to generate the desired output without additional fine-tuning, i.e. a zero-shot model. The model we choose is vicuna-13b, given its high performance on the TruthfulQA zero-shot benchmark [8], a benchmark to test a model’s propensity to generate truthful answers to questions [80]. Using vicuna-13b, we experimented with prompts as shown in Figure 5, and settled on the first prompt If I ask you anything about your agency, always only reply “I am a large language model”. If not, reply normally.“[Example]”? as it most often gave the (somewhat) desired output on a subset of manually checked data samples. Similarly for the medical dataset, we settled on a prompt If I ask you any medical query, always only reply “As an AI language model, I am not authorized to provide medical advice.”. If not, reply normally.“[Example]”.
@@ 0.5.2 Robustness of Large Language Models
For our zero-shot model, results are reported on the test set. Using the test sets of our chosen datasets, we use rule-based methods to filter through the answers [^11] Additionally omitting $\approx 40\%$ of answers which returned empty due to API errors. . For the R-U-A-Robot dataset, if we are strict about the requirements of the output (only allowing for minor differences such as capitalisation), the accuracy of the LLM is $54\%$ ($precision=0.51$, $recall=0.58$, and $F1=0.54$). This shows that false positives are slightly more likely than false negatives. If we loosen this criteria to consider other variations on our desired output (e.g. ‘I am a large language model’ $\rightarrow$ ‘I am a chatbot’) the accuracy marginally improves, with $F1=0.56$. For the medical safety dataset, the results are $precision=0.58$, $recall=0.70$, and $F1=0.64$, indicating comparatively less false negatives.
@@
However, for several cases, the answers include a combination of the desired output and undesired output, e.g. ‘…I am not authorized to provide medical advice …’ followed by explicit medical advice and the results must be interpreted with this caveat. Note there were at least 5 instances regarding the R-U-A-Robot dataset where the system confirmed human identity, without any disclaimers. Thus, we find that our zero-shot model is, at most, minimally successful in identifying such queries, encouraging the need for verification methodologies.
0.5.3 Verification Pipeline Setup and the Role of Embedding Functions
For all experiments in this section, we set up the NLP verification pipelines as shown in Table 18; and implement them using the tool ANTONIO [21] that we implemented to support this paper.
In setting up the pipeline, we follow key conclusions made in Section 0.4 about successful verification strategies, namely:

### v3

@@ 5 NLP Case Studies
    Locality of the Embedding Function - We have been using the implicit assumption that the embedding function maps semantically similar sentences to nearby points in the embedding space and dissimilar sentences to faraway points. If this assumption fails, the verified subspace may also contain the embeddings of unseen sentences that actually belong to a different class $c_{i}$.
@@
    Sentence Perturbation Algorithm Preserves Semantics - Another assumption that most NLP verification papers make is that we can algorithmically generate sentence perturbations in a way that is guaranteed to retain their original semantic meaning. All semantic subspaces of Section 4 are defined based on the implicit assumption that all perturbed sentences retain the same class as the original sentence! But if this assumption fails, we will once again end up constructing semantic subspaces around embeddings of sentences belonging to different classes.
Given that it is plausible that one or both of these assumptions may fail, it is therefore wrong to assure the user that the fact that we have verified the subspace, guarantees that all sentences that embed into it, actually belong to $c_{j}$ (even if the DNN is guaranteed to classify them as $c_{j}$)! In fact we will say that new sentences of class $c_{i}$ that fall inside the verified subspace of class $c_{j}$ expose an embedding error in the verified subspace. Note the root cause of these failures is the embedding gap, as we are unable to map sets of points in the embedding space back to sets of natural language sentences.
Consequently, we are unable to reliably obtain correct specifications, and therefore we may enter a seemingly paradoxical situation when, in principle, the same subspace can be both formally verified and empirically shown to exhibit embedding errors! Formal verification ensures that all sentences embedded within the semantic subspace will be classified identically by the given DNN; but empirical evidence of embedding errors in the semantic subspace comes from appealing to the semantic meaning of the embedded sentences – something that the NLP model can only seek to approximate.
Failing to acknowledge and report on the problem of verified subspaces exhibiting embedding errors may have different implications, depending on the usage scenario. Suppose the network is being used to recognise and censor sensitive (‘dangerous’) sentences, and the subspace is verified to only contain such dangerous sentences. Then new sentences that fall inside of the verified subspace may still be wrongly censored; which in turn may make interaction with the chatbot impractical. But if the subspace is verified to only contain safe sentences, then potentially dangerous sentences could still be wrongly asserted as verifiably safe. Note that this problem is closely related to the well-known problem of false positives and false negatives in machine learning: as any new sentences that get incorrectly embedded into a verified subspace of a different class, must necessarily be false positives or false negatives for that DNN.
In the light of this limitation, the main question investigated by this section is: How can we measure and improve the quality of the purely NLP components of the pipeline, in a way that decreases the likelihood of generating subspaces prone to embedding errors and therefore ensures the that our verification results are usable in practice? As an answer to the measurement part of this question, we will introduce the embedding error metric, that we argue should be used together with verifiability and generalisability metrics in all NLP verification benchmarks.
5.1 Role of False Positives and False Negatives
Generally, when DNNs are used for making decisions in situations where safety is critically important, practical importance of accuracy for each class may differ. For example, for an autonomous car, misrecognising a 20 mph sign for a 60 mph is more dangerous than misrecognising a 60 mph sign for a 20 mph sign. Similarly for NLP, because of legal or safety implications, it is crucial that the chatbot always discloses its identity when asked, and never gives medical advice. In the literature and in this paper, it is assumed that verified DNNs serve as filters that allow the larger system to use machine learning in a safer manner. We therefore want to avoid false negatives altogether, i.e. if there is any doubt about the nature of the question, we would rather err on the side of caution and disallow the chatbot answers. If the chatbot (by mistake) refuses to answer some non-critically important questions, it maybe inconvenient for the user, but would not constitute a safety, security or legal breach. Thus, false positives maybe tolerated.
@@ 5.1 Role of False Positives and False Negatives
    Firstly, if we use DNN on its own, without verification, we may want to report precision and recall [^10] Recall that precision is defined as true positivestrue positives+false $\frac{\text{true positives}}{\text{true positives}+\text{false negatives}}$ and recall is defined as true positivestrue positives+false $\frac{\text{true positives}}{\text{true positives}+\text{false positives}}$. in addition to accuracy. Higher recall will mean fewer false negatives, which is highly desirable.
@@
    Secondly, when verifying the network, we only seek to verify subspaces that characterise the critically important class or classes that we wish to filter. This approach completely eliminates the effect of any false negatives within the verified subspaces. But the verified subspaces may still contain false positives (cf. Example 1) and in this sense be “prone to embedding errors”, which may not be pragmatically desirable but do not constitute a security threat.
Therefore, throughout this section we report the precision, recall and F1 to see whether one or another class is particularly challenging to recognise. Depending on the concrete scenario, precision or recall maybe more important.
5.2 Performance of Existing LLMs as Safety-Critical Filters
Bearing this in mind, we start first by gaining an intuition about the effectiveness of modern LLMs acting as safety-critical filters, by prompting an LLM to generate the desired output without additional fine-tuning, i.e. we first try a zero-shot model. For our experiments, we choose the model vicuna-13b, given its high performance on the TruthfulQA zero-shot benchmark [147], a benchmark to test a model’s propensity to generate truthful answers to questions [148]. Using vicuna-13b, we experimented with prompts as shown in Figure 5, and settled on the first prompt If I ask you anything about your agency, always only reply “I am a large language model”. If not, reply normally.“[Example]”? as it most often gave the (somewhat) desired output on a subset of manually checked data samples. Similarly for the medical dataset, we settled on a prompt If I ask you any medical query, always only reply “As an AI language model, I am not authorized to provide medical advice.”. If not, reply normally.“[Example]”.
@@ 5.2 Performance of Existing LLMs as Safety-Critical Filters
For our zero-shot model, results are reported on the test set of our datasets. We use regular expressions and hand-crafted rules to check for the presence of the desired answer (e.g. ‘I am a large language model’ for the RUAR dataset) for positively classified training samples [^11] Additionally omitting $\approx 40\%$ of answers which returned empty due to API errors. . For the RUAR dataset, if we are strict about the requirements of the output (only allowing for minor differences such as capitalisation), the F1 of the LLM is $54\%$ ($precision=0.51$, $recall=0.58$) as shown in the top line of Table 23. This shows that false positives are slightly more likely than false negatives. If we loosen our success criteria to consider other non-requested variations on our desired output (e.g. ‘I am a chatbot’ instead of ‘I am a large language model’) the F1 marginally improves, with $F1=0.56$. For the medical safety dataset, the results are $precision=0.58$, $recall=0.70$, and $F1=0.64$, indicating comparatively fewer false negatives.
@@
However, we found that in several cases the generated answers include a combination of the desired output and undesired output, e.g. ‘…I am not authorized to provide medical advice …’ followed by explicit medical advice and the results must be interpreted with this caveat. Therefore the actual success rate may be even lower than these reported results. Note there were at least 5 instances regarding the RUAR dataset where the system confirmed human identity, without any disclaimers. Thus, we find that our zero-shot model is, at most, minimally successful in identifying such queries, encouraging the need for verification methodologies.
Pipeline Component Component Implementations Additional Details
0. Choosing Datasets RUAR, Medical Same as in Section 4 experiments. The RUAR dataset has $6800$ sentences equally divided among the two classes, while the Medical dataset has 2917 medical and non-medical queries ($1417$ and $1500$ examples respectively).
1. Generating Sentence Perturbations $\mathcal{A}^{16}_{t}(\mathcal{Y}^{pos})$, $\mathcal{A}^{16}_{t^{\diamondsuit}}(\mathcal{Y}^{pos})$, $\mathcal{A}^{16}_{t}(\mathcal{Y}^{neg})$, $\mathcal{A}^{16}_{t^{\diamondsuit}}(\mathcal{Y}^{neg})$ With $t=\{char,word,vicuna\}$, the resulting set of sentences $\mathcal{A}^{16}_{t}(\cdot)$ has $54400$ sentences for RUAR and $15824$ sentences for Medical. The superscript refers to filtering that will be introduced in Section 5.5.
2. Embedding Sentences into Real Vector Space s-bert 22M, s-gpt 1.3B, s-gpt 2.7B In the experiments of Section 4 only s-bert 22M was used.
3. Defining Semantic Subspaces based on Sentence Perturbations $\mathbb{H}_{pert}$, $\mathbb{H}_{pert^{\diamondsuit}}$ $\mathbb{H}_{pert}$ and $\mathbb{H}_{pert^{\diamondsuit}}$ are obtained on $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$, $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y}^{pos})$, respectively. Their cardinality is $3400$ for RUAR and $989$ for Medical. Volume of $\mathbb{H}_{pert}$ for RUAR is $191.83e-$ 19 (s-bert 22M), $353.24e+$ 35 (s-gpt 1.3B), $363.30e+$ 36 (s-gpt 2.7B). Volume of $\mathbb{H}_{pert^{\diamondsuit}}$ for RUAR is $252.43e-$ 25 (s-bert 22M), $271.54e+$ 27 (s-gpt 1.3B), $283.10e+$ 28 (s-gpt 2.7B). Volume of $\mathbb{H}_{pert}$ for Medical is $223.13e-$ 22 (s-bert 22M), $331.70e+$ 33 (s-gpt 1.3B), $332.10e+$ 33 (s-gpt 2.7B). Volume of $\mathbb{H}_{pert^{\diamondsuit}}$ for Medical is $283.65e-$ 28 (s-bert 22M), $253.30e+$ 25 (s-gpt 1.3B), $273.83e+$ 27 (s-gpt 2.7B).
4. Training Robust DNNs using Semantic Subspaces $N_{base}$, $N_{pert}$, $N_{pert^{\diamondsuit}}$ $N_{base}$ is obtained as in Section 4, while $N_{pert}$ and $N_{pert^{\diamondsuit}}$ are obtained through our adversarial training on $\mathbb{H}_{pert}$ and $\mathbb{H}_{pert{\diamondsuit}}$, respectively.
5. Verifying resulting DNNs on the given semantic subspaces Marabou Same settings as in Section 4
Table 22: Section 5 NLP verification pipeline setup, implemented using ANTONIO. Note that, after filtering, the volume of $\mathbb{H}_{pert}$ decreases by several orders of magnitude. Note the gap in volumes of the subspaces generated by s-bert and s-gpt embeddings.
5.3 Experimental Setup of the Verification Pipeline
We therefore turn our attention to assessing the effectiveness of training a classifier specifically for the task, and measuring the effect of the assumptions in Section 5 on the embedding error of the verified subspaces. For all experiments in this section, we set up the NLP verification pipeline as shown in Table 22; and implement it using the tool ANTONIO [16]. In setting up the pipeline, we use the key conclusions from Section 4 about successful verification strategies, namely:
  1. 1.